OpenStudy (anonymous):

Identify the GCF of 10x4y3 - 5x3y2 + 20x2y. 10x2y 10x3y 5x2 5x2y @rational @magepker728 @jim_thompson5910

OpenStudy (welshfella):

Firstly what is the GCF of the numbers 10 , 5 and 20 ?

OpenStudy (welshfella):

- the greatest number which will divide into all these 3 numbers

OpenStudy (anonymous):

5

OpenStudy (anonymous):

20 i mean

OpenStudy (anonymous):

nah 5

OpenStudy (welshfella):

no 20 won't divide into 5! you were right first time...

OpenStudy (anonymous):

oh okay.

OpenStudy (welshfella):

so you can eliminate a and b options already

OpenStudy (welshfella):

now what about GCF of x^2, x^3 and x^4?

OpenStudy (anonymous):

its D.

OpenStudy (welshfella):

recall the laws of indices for division x^5 / x^3 = x^(5-3) = x^2

OpenStudy (welshfella):

right

OpenStudy (anonymous):

D?

OpenStudy (welshfella):

yes